OpenStudy (anonymous):

How would I write (-4,3),(-7,-7) in standard form

OpenStudy (raden):

first, u can use the formula : (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (anonymous):

thnk you, how would I write the equation in Ax+By=C?

OpenStudy (anonymous):

in that form

OpenStudy (raden):

have u got that equation ... show to me

OpenStudy (anonymous):

I came up with y=-10/3X+16 1/3. is this correct?

OpenStudy (anonymous):

I'm being asked to put (-4,3), (-7,-7) and write it into standard form, having a hard time and was hoping that I finally figured it out

OpenStudy (raden):

lol, well look this